OpenStudy (anonymous):

Solve the equation 2/5 (x − 4) = 2x for x.

OpenStudy (anonymous):

-8 -1/2 1 -1 are my options

OpenStudy (kymber):

Cool

OpenStudy (anonymous):

x=-1 :P

OpenStudy (kymber):

o_O

OpenStudy (diyadiya):

2/5(x-4)=2x Multiply both sides by 5 2(x-4)=10x distribute 2x-8=10x Subtract 2x from both sides -8=10x-2x -8=8x x= ?

OpenStudy (kymber):

@kreshnik I like how earlier you were like 'it'd be better to give hints' and now you just gave the answer haha :p

OpenStudy (anonymous):

funny how u reported me. ur pretty intimidating i gotta say. i cried when you did that

OpenStudy (anonymous):

Ok, later @sammarino91

OpenStudy (kymber):

If you're talking to me, I didn't report you.

OpenStudy (anonymous):

@kymber I agree... But as much as I got , he's just filling an online test, and hints won't work for him! ... he doesn't even shows what he tries, doesn't even ask us how to do it he just posts it and says DO IT or SOLVE FOR X .... that's why , giving BLINDLY answer would work better for him. OR leave him aside, don't help him anymore... it's not that I want to race with anyone... not at all.. but when I see you drawing and trying to explain how it goes to him... I just feel bad, because I know he doesn't care... so just use calculator and throw him away ! LOL ;) no offense . you're doing a really good job @kymber

OpenStudy (kymber):

Yeaah, it's sad really, that he doesn't caare. :( Thaank you @kreshnik :]

OpenStudy (anonymous):

you're welcome